OpenStudy (callisto):

Sequence question #1

OpenStudy (callisto):

Before being crazy, I still have to work for a while :)

OpenStudy (anonymous):

what should i do..

OpenStudy (callisto):

Hmm... how would you do this question?

OpenStudy (callisto):

o.o ...I've work out the answer ...

OpenStudy (anonymous):

for me this will take long to figure out..

OpenStudy (anonymous):

this is an advanced topic..

OpenStudy (anonymous):

@Callisto is like a genius. She figures stuff out so quickly

OpenStudy (anonymous):

I got to admit that is true.. :)

OpenStudy (experimentx):

No sol.

OpenStudy (anonymous):

Since you work out the problem can you show the work? i'm trying to figure this out...(or is it called cheating?) @castillo

OpenStudy (callisto):

let a and l be the first and the last term Sum = (a+l)n/2 = n^3 +3n (a+l)/2 = n+3 Consider the 'sum' of the first term (a+a)/2 = 1+3 a =4 For the sum of the first 7 terms , where l= the 7th term now (a+l)/2 = n+3 (4+l)/2 = 7+3 4+l =20 l =16 So the 7th term is 16

OpenStudy (callisto):

@Open2study I wish I were a genius. But unfortunately, I'm not :(

OpenStudy (experimentx):

Oo .. i might have made some error.

OpenStudy (callisto):

Does anyone agree with my answer? I need to make sure I'm going to the right way

OpenStudy (anonymous):

I turned out to have the same answer as you did..I guess its correct? @Castillo

OpenStudy (experimentx):

i agree ... I made subtraction error

OpenStudy (experimentx):

(8^2+3*8) - (7^2+3*7) = must give you value.

OpenStudy (experimentx):

looks like I got wrong ... 4,6,8,10,12,14,16,18,20,..

OpenStudy (callisto):

(8^2+3*8) - (7^2+3*7) .... is it the 8th term?

OpenStudy (experimentx):

This must give 7th term: S7 - S6 = (7^2+3*7) - (6^2+3*6)

OpenStudy (anonymous):

Its a good thing to look back on mistakes and fixing the error.. :) It makes a good understanding..

OpenStudy (callisto):

@experimentX You're a genius!!!!!

OpenStudy (experimentx):

calculational error !!!

OpenStudy (experimentx):

lucky guess!!!

OpenStudy (callisto):

How silly am I..... :(

OpenStudy (experimentx):

right on third guess. I am glad that I'm not a doctor.

OpenStudy (anonymous):

but you are still a genius...even castillo... :) @ experimentx and Castillo

OpenStudy (experimentx):

I'm just happy to go lucky to be honest.

OpenStudy (callisto):

@experimentX Modesty :)

OpenStudy (experimentx):

haha ... a little work, a little trick, and a lots of luck.

OpenStudy (experimentx):

you are yourself pretty smart and sharp. you understand everything on a hint.

OpenStudy (callisto):

@experimentX Nope, I'm slow and silly :(

OpenStudy (experimentx):

and you manage to solve many before I do.

OpenStudy (callisto):

That's because I focus on one question, but you're looking at different question at the same time lol

OpenStudy (experimentx):

you really think that i must be super clever .... hahhah i do that because I'm out of clue ... and most of questions posted here are childish.

OpenStudy (callisto):

super clever is not enough to describe you... :P